# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ 3 Problem Description
In an RSS, three zones are deployed from center to periphery in the field (Wang et al.,, 2021): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2. Parcels are loaded to robots by workers in the loading zone. Robots drop off each parcel to its target outlet in the sorting zone to fulfill a sorting task. Waiting zone is the space where robots idle and queue behind loading stations. The area of every zone is determined in the design phase, and remains unchanged regardless of sorting demands, number of robots and workers. Let $W_{l}$ and $W_{w}$ denote the width of loading zone and waiting zone, respectively, which are considered to be constant in this study. The decision of site planning mainly focuses on the design of sorting zone, particularly the configuration of the aisle network.
@@
  • The parcels arrive in batch manner. When a parcel is loaded on a robot, the next parcel will arrive at the loading station immediately. That is, the loading time is negligible, and this parameter is considered implicitly in the loading rate of worker.
  • The arrival rate of parcels destinating for each outlet at each loading station is the same (When the sorting demand is unevenly distributed, this assumption can be achieved by increasing or decreasing the number of outlets corresponding to each destination. Additionally, outlets with extremely low and high arrival rates can be placed together to balance the flow);
  • The time for robots dropping parcels into outlets is negligible (To satisfy this assumption, robots need to release goods in advance based on its speed before reaching the target outlet, similar to the behavior of conveyor-based sorters. If it is difficult to achieve due to technical and cost issues, a 2-second drop-off time per delivery must be introduced (Zou et al.,, 2021));
  • The drop-off operation by robots around one outlet does not affect each other, i.e., each outlet allows at most four robots releasing parcels simultaneously;
  • Each robot can only carry one parcel during one delivery process;
  • Loading stations are evenly distributed in the loading zone;
  • The transportation of parcels after being collected in the outlet is not considered in this research.
  • The battery charging needs of robots are not considered.
@@
Figure 3: Illustration of aisle network in sorting zone
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term “aisle” instead of “road” in the following content to clarify that each road in the network is single-lane. A sample grid-based map in RSS is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Each cell in the map has a side length of $D$. Robots locate themselves using QR codes at the center of each cell and can only move between adjacent cells. We assume that the length of sorting robots is less than $D$, allowing them to rotate within a cell. We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots. Outlets are evenly distributed throughout the sorting zone, divided by aisles. This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density. Each loading station in the loading zone is connected to a pair of dedicated entrance and exit, which are adjacent to each other. Therefore, the number of loading stations cannot exceed the the number of aisles. Figure 3(b) shows the graph representation of the sample layout in Figure 3(a). In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), unloading node (red hexagon), or entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes. Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
@@
The sorting robots share the same kinematic parameters, with a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path finding. Common conflicts, as illustrated in Figure 4, include the following four types (Stern et al.,, 2019): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time. It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy a vertex that was occupied by another robot in the previous timestep. The latter is stationary or moving at a lower speed; (d) Deadlock conflict occurs among multiple robots, with each robot having a following conflict with another robot. Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
@@
Planing problems related to RSS can be divided into two stages: the site planning stage and operations stage. In the site planning stage, warehouse manager concentrates on the division of zones for different functions in the warehouse. The area of an RSS should be large enough to fulfill the designated throughput level in peak season, and the number of outlets should be no fewer than the number of individual sorting destinations. The decision variables of site planning stage are the number of horizontal and vertical aisles $n_{h},n_{v}$ and the number of loading stations $n_{l}$; in the operations stage, warehouse manager should adaptively adjust the number of workers and robots to minimize operations costs, given expected sorting demands. Since a warehouse is usually rented for years, it is advisable to keep the zone division unchanged during its operation. The decision variables of operations stage are the number of workers and robots in operating period $\sigma$, noted as $n_{w}^{\sigma},n_{r}^{\sigma}$, respectively.
The goal of RSS planning problem is to minimize the total costs, which consists of two parts: the facility costs $C_{f}$, and the operations costs $C_{o}$. The size of the site holds a trade-off between $C_{f}$ and $C_{o}$: A small network leads to lower capacity, thereby increasing the waiting time of robots and raising the rental cost of robots in operations costs; whereas, an overly large network leads to increased site costs and longer travel distance. An overview of research methodology is shown in Figure 5. In sections 4-6, we present the traffic management of robots, conduct system efficiency analysis, and validate the performance and accuracy of estimation formula, respectively. In section 7, a layout design optimization model is formulated to minimize the total costs of RSS given the demand.
@@ 4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput performance with the computational load of algorithms under unpredictable traffic conditions. To address multi-robot path finding and eliminate traffic gridlock within an admissible computational duration, this study adopts an innovative autonomous vehicle management scheme, called Rhythmic Control (RC) (Lin et al.,, 2021). It enables the planning of uninterrupted scheduling for robot fleets. In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle. By leveraging the reservation mechanism, we propose a new centralized framework called Rhythmic Control in Sorting Scenario (RC-S). RC-S presents high orderliness and serves as the foundation for theoretical analysis of system efficiency, which will be discussed in detail in section 5.
@@ 4.1 Virtual platoon and cycle
Lin et al., (2021) investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). VPs represent the spatiotemporal slot that is generated in a rhythmic manner and keep uniform linear motion. As shown in Figure 6(a), VPs move along a designated spatiotemporal trajectory and maintain a constant safe distance between each other. Figure 6(b) shows the spatiotemporal trajectory of VPs in a four-lane network, without any collision or deceleration. We discretize time based on the generation time of VPs at the entrance, and each time interval is noted as a cycle. The duration of each time interval is the cycle length. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass one intersection alternately with relatively short headway and the intersection capacity could be maximized.
Whenever a vehicle needs to pass through a lane, it is mandated to follow the movement of one VP in that lane until it reaches its destination and exits the lane. Before entering the network, the vehicle must wait for an unoccupied VP at the entrance. It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity from the corresponding VPs in different segments of the route. The capacity is released when the vehicle leaves each VP. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Consequently, the trajectories of vehicles are conflict-free.
@@
(b) Trajectory of virtual platoons in different directions. (Lin et al.,, 2021)
@@
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control. Within this framework, it is essential to clarify the movement patterns of each entity for integration into a sorting-specific robot management system. For detailed information on the RC framework, we refer readers to Lin et al., (2021). The following points provide a foundational overview critical to our study on the implementation of the RC-S scheme:
  • A VP represents a time–space slot that is available for being occupied by vehicles and is generated in a rhythmic manner in each entrance of the network.
  • VPs move in a straight line at a constant speed while maintaining a fixed distance, ensuring that the total number of VPs present in the road network remains the same at any given moment.
  • After entering the network, vehicles need to continuously move along with a specific VP or transfer from one VP to another until exiting through an exit.
  • At any given moment, the number of vehicles accommodated by a VP cannot exceed its capacity. If the required VP capacity for vehicle travel is insufficient, the vehicle needs to queue and wait at the entrance until entering the network in a subsequent cycle.
@@ 4.2 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver. To focus on the behavior of each individual robot, we discretize the map into grids, defining each grid where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{e}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance/exit nodes and edges, respectively. To reduce intersections, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Based on the network structure, each VP can accommodate at most one AGV. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each unit cell in the grid map and the cycle length of rhythmic control, respectively. Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning. Robots enter the network with fixed interval $\tau_{c}$, thus the upper bound of robot flow on each grid is $1/\tau_{c}$.
@@
In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. We have:
@@
$\tau_{e}$ and $\tau_{c}$ are determined by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
@@
where $D$ is the length of each grid. At phase(a) of a cycle, new VPs are released at the entrance of each aisle; at phase(c), VPs leave sorting zone through the exit of each aisle. Thus, if we do not count the VPs at the exit of each aisle, then the number of VPs in the network remains the same throughout the entire sorting process. Intersections are occupied by VPs at phases(a) and (c); each outlet could be served by four VPs in four different directions at phases(b) or (d). The fixed number of VPs and the regularity in their phases allow us to manage and analyze this system from a global perspective. The speed of each VP could be derived by:
@@
Figure 8: Illustration of turning and transfer process for one robot (red dot)
RSS does not require auxiliary turning lanes since robots are capable to rotate within the intersection grid. The process of turning and transfer between VPs for robots are demonstrated in Figure 8. As shown in the figure, the robot to turn first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP and follows the movement of it. Turning action entails one cycle for both the original and subsequent VPs, leading to a reduction in traffic capacity.
During this process, the parameters of RC-S needs to satisfy the following conditions:
@@
Constraint (5) implies that the robot can catch up with the VP in a turn, where $d(\tau_{e},v_{VP},v_{max},c_{max})$ denotes the maximum travel distance a robot can cover in acceleration/deceleration process. The expression is derived in detail in Appendix B; Constraint (6) imposes limits on the rotation speed $\omega_{r}$, ensuring that the robot can complete a 90-degree turn within one phase. In summary, using expressions (1)-(6) and the kinematic parameters of the robots, we could determine the parameters for RC-S.
The trajectory of robot in one delivery task should start from one loading station, pass the target outlet corresponding to the destination of parcel, and end at another loading station. We define feasible path in RC-S:
@@
A feasible path in RC-S is a sequence of VPs and their occupying cycles. The trajectories traveled by these VPs during the occupied cycles are connected end-to-end, linking two active loading stations at the periphery of the network.
@@
Definition 1 defines the candidate spatio-temporal paths for robots under RC-S, where an active loading station indicates that a worker is currently engaged in tasks. When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, $i,j\in\mathcal{L}$, where $\mathcal{L}$ is the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. Then, we denote the set of all feasible path as $\mathcal{R}=\cup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing the paths start from loading station $i$, pass by outlet $k$ and end at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing by outlet $k$, and ^$\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths heading for station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and ^$\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$. In each cycle, the feasible path assignment problem (FPA) for all the loaded robots is formulated as follows.
$\underset{x_{i,k}^{r},\hat{x}_{i,k}}{\min}\ \sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}(\sum_{r\in\mathcal{R}_{i,k}}c_{i,k}^{r}x_{i,k}^{r}+\hat{c}_{i,k}\hat{x}_{i,k})$
$s.t.\quad$ $\sum_{i\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}\delta_{i,k}^{r,\nu,l}x_{i,k}^{r}\leq N_{\nu}^{l}$ $\forall\nu\in\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{e},l\in\mathcal{C}$ (7)
$\hat{x}_{i,k}+\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=d_{i,k}$ $\forall i\in\mathcal{L},k\in\mathcal{O}$ (8)
$\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{i,k}}x_{i,k}^{r}=\sum_{j\in\mathcal{L}}\sum_{k\in\mathcal{O}}\sum_{r\in\mathcal{R}_{j,k}\cap\hat{\mathcal{R}}_{i}}x_{j,k}^{r}$ $\forall i\in\mathcal{L}$ (9)
$x_{i,k}^{r},\hat{x}_{i,k}\in\{0,1\}$ $\forall i\in\mathcal{L},k\in\mathcal{O},r\in\mathcal{R}_{i,k}$ (10)
There are two sets of variables: $x_{i,k}^{r}$ equals one if the feasible path $r\in\mathcal{R}_{i,k}$ is reserved in current cycle by a robot, and zero otherwise; $\hat{x}_{i,k}$ equals one if the robot at loading station $i$ is heading for outlet $k$ and is kept waiting until the next cycle, and zero otherwise. In the objective function, $c_{i,k}^{r}$ is the travel time on path $r\in\mathcal{R}_{i,k}$, and $\hat{c}_{i,k}$ is the penalty of delay. The objective function is to minimize the total costs in this cycle. Constraint (7) indicates that in each cycle, at most one robot can be present at node $\nu$, where $N_{\nu}^{l}$ is the updated remaining capacity in node $\nu$ in cycle $l$, $N_{\nu}^{l}\leq 1$. Parameter $\delta_{i,k}^{r,\nu,l}$ equals one if feasible path $r$ will occupy node $\nu$ in cycle $l$ and zero otherwise. Constraint (8) describes the path allocation results for a robot at loading station $i$, where $d_{i,k}$ equals one if the robot is heading for outlet $k$, and zero otherwise. Constraint (9) is the conservation of flow constraint, ensuring a stable number of robots at each loading station.
In large-scale network, the number of feasible paths rapidly increases. It is difficult to solve FPA within a reasonable amount of time. Here we propose a heuristic algorithm to generate a feasible solution in a relatively short time, as shown in Algorithm 1. To expand the search depth, feasible paths starting at some future cycle are allowed to be reserved. We further define the entry cycle:

### v4

@@ 3 Problem Description
In an RSS, three zones are deployed from center to the periphery of the field (10.1007/978-981-16-8174-5_10): 1) sorting zone; 2) loading zone; 3) waiting zone, as shown in Figure 2. Parcels are loaded to robots by workers in the loading zone. Robots drop off each parcel at its target outlet in the sorting zone to fulfill a sorting task. The waiting zone is the space where robots idle and queue behind loading stations. The area of every zone is determined during the design phase and remains unchanged regardless of sorting demands or the numbers of robots and workers. Let $W_{l}$ and $W_{w}$ denote the widths of the loading zone and the waiting zone, respectively, both considered constant in this study. The decision of site planning mainly concerns the configuration of the aisle network in the sorting zone.
@@
  • The parcels arrive in batches. When a parcel is loaded onto a robot, the next parcel will arrive at the loading station immediately. That is, the loading time is negligible and is implicitly captured in the loading rate of the workers.
  • The arrival rate of parcels destined for each outlet at each loading station is identical. (When sorting demand is unevenly distributed, this can be achieved by adjusting the number of outlets corresponding to each destination. Additionally, outlets with extremely low or high arrival rates may be grouped together to balance the flow.)
  • Robots release parcels without slowing down or stopping, and drop-off time is considered negligible. (To satisfy this assumption, robots must release parcels in advance based on their speed before reaching the target outlet, similar to the behavior of conveyor-based sorters.);
  • Drop-off operations around an outlet are independent, i.e., each outlet allows at most four robots releasing parcels simultaneously.
  • Each robot can carry only one parcel during one delivery process;
  • Once a robot exits the network through an exit node, it immediately joins the queue of the associated loading station without repositioning between loading stations. Travel time within the waiting and loading zones is neglected.
  • The transportation of parcels after they are collected at the outlet is not considered.
  • Battery charging requirements of robots are not considered.
@@
Figure 3: Illustration of aisle network in the sorting zone
To better evaluate the performance of this system with high loads and complicated sorting demands, the microscopic traffic management of robots in the sorting zone needs careful investigation. Notably, we use the term “aisle” instead of “road” in the following content to clarify that each road in the network is single-lane. A sample grid-based map in RSS is depicted in Figure 3(a). It consists of $n_{v}$ vertical aisles and $n_{h}$ horizontal aisles. Each cell in the map has a side length of $D$. Robots locate themselves using QR codes at the center of each cell and can only move between adjacent cells. We assume that the length of sorting robots is less than $D$, allowing them to rotate within a cell. We use the term “outlet” to refer to the special cells designated for collecting parcels delivered by robots. Outlets are evenly distributed throughout the sorting zone, seperated by aisles. This layout allows robots to deliver parcels to outlets on both sides of the aisle and ensures that the outlets are arranged with sufficient density.
Each loading station in the loading zone is connected to a pair of adjacent entrance and exit nodes. Therefore, the number of loading stations cannot exceed the number of aisles. A dedicated worker is required at each loading station to perform parcel loading operations. A loading station is defined as active if a worker is assigned and engaged in tasks at that station. Figure 3(b) shows the graph representation of the sample layout in Figure 3(a). In the graph, each cell that makes up the aisles is modeled as either a conflict node (blue dot), an unloading node (red hexagon), or an entrance/exit node (green/yellow square), with edges of length $D$ connecting adjacent nodes. Conflict nodes are at the intersections of aisles, unloading nodes are where robots deliver parcels, and entrance/exit nodes are at aisle ends.
@@
The sorting robots share the same kinematic parameters, with a maximum velocity of $v_{max}$ and maximum acceleration/deceleration of $c_{max}$. Robots possess the capability of in-place rotation, with angular velocity denoted as $\omega_{r}$. To ensure collision-free movement of robots, the robot management system needs to avoid conflicts in path planning. Common conflicts, as illustrated in Figure 4, include the following four types (stern2019multi): (a) Swapping conflict occurs when two robots are planned to swap locations at the same time. It does not exist in a one-way network; (b) Vertex conflict occurs when more than two robots are planned to occupy the same vertex; (c) Following conflict occurs when one robot is planned to occupy the vertex currently occupied by another robot, while the latter is either stationary or moving at a lower speed; (d) Deadlock conflict involves multiple robots, each trying to maintain the minimum headway with another, resulting in mutual blockage. Among these conflicts, both swapping conflicts and deadlock conflicts necessitate a re-planning of paths. The occurrence of either can result in significant losses in system efficiency.
In this study, the RSS layout planning problem is structured as a two-stage cost-minimization model subject to throughput constraints, integrating long-term layout design with operational-level traffic control of robots:
  1. 1.
    Site planning stage (strategic level): The objective is to minimize the total cost, including facility cost $C_{f}$ and expected operations cost $C_{o}$ across different operating periods $\sigma\in\mathcal{S}$:
    $\min\ C_{d}=C_{f}+C_{o}$ (1)
    Each period $\sigma$ is associated with a required throughput $T^{\sigma}$, representing changes in sorting demand. Decision variables include the number of horizontal aisles $n_{h}$, vertical aisles $n_{v}$, loading stations $n_{l}$, and the number of workers $n_{w}^{\sigma}$ and robots $n_{r}^{\sigma}$ allocated in each period $\sigma$. Layout and resource configurations must meet the throughput requirements:
    where $\tilde{T}_{O}$ is the estimated system throughput. A trade-off exists between $C_{f}$ and $C_{o}$: larger networks increase facility costs and travel distances, while smaller layouts may cause congestion and reduce network capacity.
  2. 2.
    Operations stage (operational level): Given a fixed layout and resource configurations, the goal is to coordinate robot traffic efficiently to achieve stable and predictable system throughput. The robot traffic management method regulates robot movements and trajectory assignments, directly determining the system throughput. The estimation of system throughput $\tilde{T}_{O}$ serves as a critical link between operational performance and upper-level planning.
@@
Figure 5 provides an overview of this integrated framework. In Sections 46, we introduce the robot traffic management approach, analyze the system efficiency, and validate the performance and accuracy of the proposed throughput estimation method. In Section 7, we formulate the layout optimization model to minimize the total cost of RSS under given sorting demand.
@@ 4 Management of High-density Robot Traffic in RSS
Managing a large number of robots within a dense network poses significant challenges, especially when balancing system throughput with the computational load of algorithms under unpredictable traffic conditions. To address the multi-robot path finding problem and eliminate traffic gridlock within an acceptable computational time, this study adopts an innovative autonomous vehicle management scheme called Rhythmic Control (RC) (RCII), which enables uninterrupted scheduling of robot fleets. In this section, we begin with the key concepts inherited from the original RC: the virtual platoon and the cycle. By leveraging the reservation mechanism of RC, we propose a new centralized framework called Rhythmic Control for the Sorting Scenario (RC-S). RC-S exhibits a high degree of orderliness and serves as the foundation for the theoretical analysis of system efficiency, which will be discussed in detail in Section 5.
@@ 4.1 Virtual platoon and cycle
RCII investigated a method of incorporating rhythm into traffic management at the macro-traffic level and introduced the concept of virtual platoon (VP). A VP is a reserved spatio-temporal slot along a lane, characterized by a uniform linear trajectory and generated at fixed intervals. Each VP could be identified by its corresponding lane and the interval in which it is released. As shown in Figure 6(a), VPs move along a designated spatio-temporal trajectory and maintain a constant safe distance between each other. Figure 6(b) shows the spatio-temporal trajectory of VPs in a four-lane network, without any collision or deceleration. We discretize time based on the generation time of VPs at the entrance, and each time interval is referred to as a cycle. The duration of each time interval is the cycle length. By staggering the entry times of VPs at different entrances within the same cycle, VPs pass a intersection alternately with relatively short headway and the intersection capacity could be maximized.
Whenever a vehicle needs to pass through a lane, it is required to follow the movement of a VP in that lane until it reaches its destination and exits the lane. Before entering the network, the vehicle must wait for an unoccupied VP at the entrance. It first sends its destination to the control system, which then plans a conflict-free route and reserves one unit of capacity in the corresponding VPs in different segments of the route. The capacity is released when the vehicle leaves each VP. Through the reservation mechanism, RC ensures that the number of vehicles within each VP does not exceed its predetermined capacity. Consequently, the trajectories of vehicles are conflict-free.
@@
(b) Trajectory of virtual platoons in different directions. (RCII)
@@
RC is essentially a traffic control method, similar to a combination of globally coordinated traffic signal control and road network flow control. Within this framework, it is essential to clarify the movement patterns of each entity so that RC can be integrated into a sorting-specific robot management system. For detailed information on the RC framework, we refer the reader to RCII. The following points provide a foundational overview that is critical to our study on the implementation of the RC-S scheme:
  • A virtual platoon (VP) represents a spatio-temporal slot that can be reserved and occupied by robots and is generated in a rhythmic manner at each entrance of the network.
  • VPs move in a straight line at a constant speed while maintaining a fixed spacing, ensuring that the total number of VPs present in the aisle network remains constant.
  • After entering the network, a robot must continuously follow a specific VP or transition to another VP until it exits.
  • The capacity of each VP is one, which means at most one robot can occupy it at any time.
  • A robot may enter the network only when the required VP capacity is available; otherwise, it must wait for a later cycle.
@@ 4.2 Concept of RC-S
In order to embody RC in the RSS, we regulate the behavior of robots, and propose the RC-S scheme, which can be regarded as a centralized MAPF solver. Under RC-S, the behaviors of robots are limited to three actions: moving straight, dropping off parcel, and turning. We discretize the map into grid-cells, defining each grid-cell where robots can pass through as a node. Given $(n_{v},n_{h})$, the network in the sorting zone is represented by a directed graph $\mathcal{G}=(\mathcal{V}=\mathcal{V}_{c}\cup\mathcal{V}_{u}\cup\mathcal{V}_{en}\cup\mathcal{V}_{ex},\mathcal{E})$, where $\mathcal{V}_{c}$, $\mathcal{V}_{u}$, $\mathcal{V}_{en}$, $\mathcal{V}_{ex}$ and $\mathcal{E}$ stand for the set of conflict nodes, unloading nodes, entrance nodes, exit nodes and edges, respectively. To reduce potential conflicts and increase the density of outlets, we adopt a one-way network, where aisles from different directions are arranged in an interleaved pattern, as shown in Figure 7. Based on the network structure, each VP can accommodate at most one robot. Let $\tau_{e}$ and $\tau_{c}$ denote the fixed travel time of VPs on each grid-cell and the cycle length of rhythmic control, respectively. That is, at each entrance node, VPs enter the network at fixed intervals of $\tau_{c}$.
@@
In each cycle, the distribution of VPs could be divided into four phases (Figure 7(a)-(d)), and the duration of each phase is $\tau_{e}$. Each VP moves forward one grid-cell along its assigned direction after $\tau_{e}$, transitioning the system from one distribution phase to the next, in the sequence: (a) $\rightarrow$ (b) $\rightarrow$ (c) $\rightarrow$ (d) $\rightarrow$ (a) $\rightarrow\dots$. We could obtain:
@@
$\tau_{e}$ and $\tau_{c}$ are restricted by the maximum velocity of robot $v_{max}$ and the maximum loading rate of each loading station $r_{l}$, respectively:
@@
where $D$ is the length of each grid-cell. At phase(a) of a cycle, new VPs are released at the entrance node of each aisle; at phase(c), VPs leave the sorting zone through the exit node of each aisle. Thus, if we do not count the VPs at the exit nodes, then the number of VPs in the network remains the same throughout the entire sorting process. Each intersection is occupied by two different VPs at phases (a) and (c) within a cycle, respectively. Each outlet could be served by four VPs in four different directions at phases(b) or (d). The fixed number of VPs and the regularity in their phases allow us to manage and analyze this system from a global perspective. The speed of each VP could be derived by:
@@
Figure 8: Illustration of turning and transfer process for a robot
Figure 9: (a) Intersection (b) Trajectory of a turning robot
RSS does not require auxiliary turning lanes, since robots are able to rotate within the intersection grid-cell. The process of robot turning and transfer between VPs is demonstrated in Figure 8. As shown in the figure, the turning robot first decelerates to zero, stops following its original VP, rotates 90 degrees at the intersection, joins a new VP, and then follows its movement. The turning action occupies one cycle in both the original and the subsequent VPs, leading to a reduction in traffic capacity.
During this process, the parameters of RC-S must satisfy the following conditions:
@@
Constraint (7) imposes a restriction on $\tau_{e}$ based on the robot’s maximum acceleration/deceleration rate $c_{max}$. The function $d(\tau_{e},v_{VP},v_{max},c_{max})$ represents the maximum distance that a robot can cover during acceleration and deceleration, and its derivation is provided in Appendix B. Constraint (8) sets a minimum requirement on $\tau_{e}$ to ensure that the robot can complete a 90-degree turn within one phase, given its rotation speed $\omega_{r}$. As a result, using Expressions (3)–(8) together with the kinematic parameters of the robots, we could determine the parameters for RC-S.
The trajectory of a robot for a single delivery task starts at a loading station, passes the target outlet corresponding to the parcel’s destination, and ends at another loading station. We define a feasible path under the RC-S scheme as follows:
@@
A feasible path in RC-S is a sequence of VP–cycle pairs, denoted by $\{(p_{1},l_{1}),(p_{2},l_{2}),\ldots,(p_{n},l_{n})\}$, where $(p_{i},l_{i})$ indicates that the robot occupies VP $p_{i}$ during cycle $l_{i}$. The sequence must satisfy the following conditions:
  1. 1.
    Start and end points: The first VP $p_{1}$ originates from an entrance node $\nu_{en}\in\mathcal{V}_{en}$ in cycle $l_{1}$, and the last VP $p_{n}$ ends at an exit node $\nu_{ex}\in\mathcal{V}_{ex}$ in cycle $l_{n}$.
  2. 2.
    Temporal consistency: For every consecutive pair $(p_{i},l_{i})$ and $(p_{i+1},l_{i+1})$:
    • If $p_{i+1}=p_{i}$, then $l_{i+1}=l_{i}+1$ (the robot remains in the same VP).
    • Otherwise, $l_{i+1}=l_{i}$ (the robot transitions between VPs).
  3. 3.
    Spatial connectivity: If $p_{i}\neq p_{i+1}$, then $p_{i}$ and $p_{i+1}$ must intersect at a common conflict node $\nu_{c}$ from different directions, and the corresponding entry times $t_{i}$ and $t_{i+1}$ must satisfy:
    $\tau_{c}\cdot(l_{i}-1)\leq t_{i}<t_{i+1}<\tau_{c}\cdot\left(l_{i}+\frac{1}{2}\right),\quad t_{i+1}=t_{i}+\frac{\tau_{c}}{2}$
    where $\tau_{c}$ denotes the cycle length.
PROPOSITION 1.
Given $\mathcal{G}=(\mathcal{V},\mathcal{E})$, each feasible path under the RC-S control scheme can be represented as a sequence of node–cycle pairs $\{(\nu_{i},l_{i})\}$.
The proof of Proposition 1 is provided in Appendix B. Definition 1 specifies the candidate spatio-temporal trajectory patterns for robots under RC-S. Condition 1 ensures that the trajectory connects valid start and end nodes on the network boundary. Conditions 2 and 3 guarantee continuity and compliance with RC-S by restricting robot movements in accordance with the RC-S rules. Moreover, since robots are not permitted to stop except when turning, feasible paths cannot contain transfers between VPs in the same direction. Proposition 1 enables the transformation of VP reservations into node–cycle pairs, forming the basis for the optimization model presented later.
@@
When a robot performs a sorting task, it moves from one loading station $i$ to another loading station $j$, with $i,j\in\mathcal{L}$, where $\mathcal{L}$ denotes the set of all loading stations. Let $\mathcal{O}$ represent the set of outlets. We denote by $\mathcal{R}$ the set of all feasible paths, that is, $\mathcal{R}=\bigcup_{i\in\mathcal{L},j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(i,k,j)$, where $\mathcal{R}(i,k,j)$ is the subset containing feasible paths starting from loading station $i$, passing outlet $k$, and ending at loading station $j$. To simplify the notation, we use $\mathcal{R}_{i,k}$ to represent the set of all feasible paths originating from station $i$ and passing outlet $k$, and $\hat{\mathcal{R}}_{i}$ to represent the set of all feasible paths terminating at station $i$, that is, $\mathcal{R}_{i,k}=\cup_{j\in\mathcal{L}}\mathcal{R}(i,k,j)$ and $\hat{\mathcal{R}}_{i}=\cup_{j\in\mathcal{L},k\in\mathcal{O}}\mathcal{R}(j,k,i)$.
The number of paths in $\mathcal{R}(i,k,j)$ is theoretically infinite, since without additional constraints, each feasible path could include an arbitrary number of turns. We limit the total number of feasible paths by allowing at most three turns and prohibiting robots from stopping within the network except for turning. These restrictions ensure that $\mathcal{R}$ is finite and that each feasible path uses no more than four distinct VPs. The rationale behind this choice will be explained in the analysis of the average travel distance in the next section.
To expand the search space, feasible paths starting at future cycles should also be considered. We further define the entry cycle: